# SelfDefend: LLMs Can Defend Themselves against Jailbreaking in a Practical Manner

### v1

@@ 3.1 Design Objectives
Handling all kinds of jailbreak attacks. The proposed defense should be able to handle all categories of jailbreak attacks listed in §2, including not only optimization-based adversarial jailbreaks but also template-based, indirect, and multilingual jailbreaks.
@@
Incurring negligible delay to user prompts. The proposed defense should not impact the user experience, causing either no delay or only a negligible one to normal user prompts.
@@
Providing explanations for potential jailbreak queries. Similar to O2, when the proposed defense detects any query potentially related to jailbreaking, it should provide helpful explanations on why the query is considered harmful.
@@
Compatible with both open-source and closed-source LLMs. The proposed jailbreak defense approach should protect both white-box, open-source LLMs as well as black-box, closed-source LLMs.
We clarify that compared with the other three objectives, O3 is not a mandatory objective. Nevertheless, we believe O3 is particularly important for users to forensically understand potential jailbreak queries and better protect against jailbreak attacks in the future.
@@ 3.2 Analysis of Existing Defenses
Table I summarizes our analysis of major LLM jailbreak defenses under the four objectives we envisioned above. They can be roughly categorized into tuning-based (the first six rows) and non-tuning-based (the last seven rows) mechanisms. Specifically, tuning-based defenses aim to fundamentally improve a model’s safety alignment against jailbreaking, while non-tuning-based defenses can be typically applied to any off-the-shelf LLMs to enhance their safety against jailbreak attacks. In the rest of this subsection, we present our analysis results from four perspectives:
@@
First, most jailbreak defenses target multiple kinds of jailbreak attacks but typically do not cover advanced indirect attacks (O1: ), while a few existing defense mechanisms are specifically designed to defend against optimization-based adversarial attacks only (O1: ). The latter includes two perplexity-based filtering approaches [41, 62] and several input perturbation-based approaches [49, 50, 51, 52]. Specifically, Alon and Kamfonas [41] proposed the first tuning-based defense in the sense that they not only leveraged perplexity values as an indicator to detect prompts with adversarial suffixes but also tuned a classifier to consider both the sequence length of prompts and their perplexity for improved filtering. Another line of GCG-specific jailbreak defenses perturbed copies of the input prompt and aggregated the output responses, with SmoothLLM [51] as a representative example.
@@
Second, the majority of prior defenses inherently incur additional delays to user prompts (O2: ), with only two exceptions. Since the design principles of most prior defenses are to check the target LLM’s responses [47, 48, 42] and internal statuses [41, 62, 45, 54, 55, 46], or to conduct extra-round analyses of the input prompt [49, 50, 51, 52, 53], it is thus inherently difficult for these approaches to avoid additional delay. The only exceptions are [73, 74, 43, 44], which conducted optimization-based suffix or prefix training for defensive purposes, with the training objective being the exact opposite of the GCG attack. As a result, the tuned models behave like normal LLMs, incurring no extra delay to user prompts.
@@
Third, most non-tuning-based defenses have the potential to provide explanations for jailbreak queries (O1: or ), whereas half of the tuning-based defenses cannot because they rely solely on LLMs’ internal indicators for tuning (O3: ). To provide explanations for potential jailbreak queries, a defense mechanism needs to understand the semantics of incoming prompt queries. Hence, it is hard for approaches that rely solely on LLMs’ internal indicators, whether they are tuning-based [45, 46] or non-tuning-based [54, 55], to provide straightforward explanations to users.
@@
Fourth, most non-tuning-based defenses are compatible with both open-source and closed-source LLMs (O4: ), except for those that need to leverage LLMs’ internal indicators, while the opposite is true for tuning-based defenses (O4: ). Unless they need to monitor LLMs’ internal indicators [48, 54, 55], non-tuning-based defenses, such as Self Defense [47] and IAPrompt [53], are naturally compatible with both open-source and closed-source LLMs. On the other hand, tuning-based defenses typically require white-box access to the LLMs to enable their in-depth defense. The only exception is Meta’s Llama Guard [42], which aligns the classifications of both prompt and response, and thus, theoretically, it can serve as a plug-in model to work with both open-source and closed-source LLMs.
Figure 1: A high-level overview of the SelfDefend framework and our study.
@@ 4 The SelfDefend Framework
Based on this idea, we propose a generic LLM jailbreak defense framework called SelfDefend. As shown in Figure 1, SelfDefend creatively establishes a shadow stack alongside the normal stack in the LLM space to conduct checkpoint-based access control, which mimics traditional security defense concepts such as the shadow stack for defending against buffer overflow attacks [56] and the library-based checkpoint from SCLib [75]. We denote the target LLM in the normal stack as $LLM_{target}$ and the defense LLM in the shadow stack as $LLM_{defense}$. SelfDefend simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate. In SelfDefend, $LLM_{defense}$ can be instantiated from the same model as $LLM_{target}$, although in practice we suggest using a dedicated $LLM_{defense}$ that is robust and low-cost for detecting jailbreak queries. Since $LLM_{defense}$ does not need to modify $LLM_{target}$ or know any of its internal details, SelfDefend can protect all kinds of target LLMs, whether they are open-source or closed-source (O4: ).
Given an incoming prompt query $P_{query}$, SelfDefend dispatches it to both $LLM_{target}$ and $LLM_{defense}$ for concurrent processing. Specifically, $LLM_{target}$ processes $P_{query}$ as usual, whether it is a normal user prompt or an adversarial jailbreak prompt, but caches its token-by-token output until a checkpoint is triggered from the shadow stack. In contrast, $LLM_{defense}$ employs a dedicated detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$). Once a token of “No” (indicating no issue) or a harmful portion (prompt/intention) is output, $LLM_{defense}$ triggers a checkpoint in the normal stack so that $LLM_{target}$ can decide whether to release the token-by-token output. Since the time for outputting the token “No” is very limited, the additional delay introduced by SelfDefend is negligible for normal user prompts (O2: ). Moreover, the identified harmful portion can help generate an explainable LLM response to adversarial prompts (O3: ). These unique advantages make SelfDefend the first practical jailbreak defense compared to the existing defense mechanisms explained in §3.2.
Now the effectiveness of SelfDefend depends on whether it can handle all kinds of jailbreak attacks (i.e., O1). For this, we proceed to conduct an empirical study in §5 to validate our earlier hypothesis that “LLMs have the capability to identify harmful portions in user queries.”
@@ 5 An Empirical Measurement
In this section, we conduct extensive measurements to empirically show that existing LLMs, notably the commonly used GPT-3.5/4 models, perform acceptably under the SelfDefend framework. This not only validates the hypothesis assumed in §4 but also enables widespread, convenient usage of SelfDefend without the need for further tuning of a defense model. For example, various custom GPTs applications [76] can adopt SelfDefend to immediately enhance their defense capabilities against jailbreaks.
Figure 2: SelfDefend’s two detection prompt templates.
5.1 Initializing SelfDefend for GPT-3.5/4
@@
To facilitate measurements, we need to first initialize SelfDefend for GPT-3.5/4. For GPT-3.5-based experiments, both $LLM_{target}$ in the normal stack and $LLM_{defense}$ in the shadow stack are set to GPT-3.5. Likewise, for GPT-4-based experiments, both $LLM_{target}$ and $LLM_{defense}$ are set to GPT-4.
@@
Prompt Templates. $LLM_{target}$ directly uses user prompts $P_{query}$ but we need to set up the detection prompts for $LLM_{defense}$. As mentioned earlier in §4, we have designed $P_{direct}$ and $P_{intent}$ for two different defense schemes. Figure 2 shows their detailed prompt templates. We can see that $P_{direct}$ aims to directly pinpoint the harmful part from the original jailbreak/normal $P_{query}$; otherwise, it outputs a token of “No.” In contrast, $P_{intent}$ employs a chain-of-thought (CoT) [77] approach to first summarize the true intention of $P_{query}$ and then, based on the summarized intention, further extract any safety-violation part, if present. It also includes two examples in the prompts to help LLMs better understand the instructions and conduct few-shot in-context learning [78].
By comparing these two prompts, we expect that $P_{direct}$ would perform better on optimization-based jailbreaks that directly include the jailbreak portion, while $P_{intent}$ might be more effective against indirect jailbreaks. One additional objective of the empirical measurement in this section is to explore the best setup for deploying SelfDefend with off-the-shelf LLMs under four different combinations (GPT-3.5 vs. GPT-4; $P_{direct}$ vs. $P_{intent}$).
Environment Setup. In our implementation, we use the OpenAI API in Python to access GPT-3.5 and GPT-4. The API versions of GPT-3.5 and GPT-4 are gpt-3.5-turbo-0125 and gpt-4-0125-preview, respectively. We set the hyper-parameters max_tokens, temperature, n, and top_p to 128, 0, 1, and 0, respectively. When accessing the GPT APIs, no system prompt is added.
TABLE II: The details of our collected benchmark datasets.
@@
JailbreakBench [79] 100 GCG [29], JSAA [31]
PAIR [36], AutoDAN [28]
@@
TABLE III: The ASR results from testing LLMs against four major categories of jailbreak attacks and normal prompts.
LLMs Optimization-based Template-based Indirect Multilingual Normal
GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 (baseline) 0.56 0.94 0.72 0.90 0.78 0.98 0.39 0.97
GPT-3.5-based Shadow Stack ($P_{direct}$) 0.72 0.98 0.77 0.96 1.00 0.98 0.87 0.97
GPT-3.5-based SelfDefend ($P_{direct}$) 0.45 0.92 0.60 0.87 0.78 0.96 0.36 0.95
GPT-3.5-based Shadow Stack ($P_{intent}$) 0.28 0.06 0.37 0.35 0.86 0.22 0.52 0.99
GPT-3.5-based SelfDefend ($P_{intent}$) 0.19 0.04 0.29 0.31 0.71 0.22 0.20 0.97
Reduction factor 66.07% 95.74% 59.72% 65.55% 8.97% 77.55% 48.71% 0%
GPT-4 (baseline) 0.08 0.16 0.33 0.19 0.74 0.90 0.08 0.97
GPT-4-based Shadow Stack ($P_{direct}$) 0.01 0.00 0.11 0.01 0.05 0.27 0.14 0.96
GPT-4-based SelfDefend ($P_{direct}$) 0.00 0.00 0.10 0.01 0.04 0.26 0.01 0.94
Reduction factor 100% 100% 36.36% 94.73% 94.59% 71.11% 87.50% 3.09%
GPT-4-based Shadow Stack ($P_{intent}$) 0.08 0.05 0.21 0.07 0.13 0.36 0.30 0.99
GPT-4-based SelfDefend ($P_{intent}$) 0.05 0.05 0.18 0.07 0.13 0.28 0.01 0.97
@@
Figure 3: The extra delay $\Delta d$ caused by SelfDefend for normal prompts and jailbreak prompts, respectively.
5.2 Datasets and Attack Setup
Benchmarks. Based on the four categories of existing jailbreak attacks we surveyed in §2 — optimization-based, template-based, indirect, and multilingual jailbreaks — we identify representative jailbreak attack methods in each category. We then collect three benchmark datasets, JailbreakBench [79], MultiJail [34], and AlpacaEval [57], from which we use their user prompts for testing SelfDefend. Table II lists the details of our collected benchmark datasets. Specifically, we use a set of 100 instructions from JailbreakBench [79], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [29] and JSAA [31]), template-based jailbreaks (PAIR [36] and AutoDAN [28]), and indirect jailbreak attacks (DrAttack [32] and Puzzler [33]). In contrast, we directly use the original prompts from the MultiJail and AlpacaEval datasets to construct inputs for the scenarios of multilingual jailbreaks and normal prompts, respectively.
Attack Setup. For GCG, we choose its individual version and optimize the suffix on Vicuna-7b-v1.3 [80] with a batch size of 512 and 500 optimization steps. For JSAA, we query GPT-3.5 with 10,000 iterations and 10 restarts. We perform transfer attacks on GPT-4 using the prompts generated on GPT-3.5. For PAIR, we select Vicuna-13b-v1.5 [80] as the attack model. For AutoDAN, we choose its first version of the genetic algorithm, i.e., AutoDAN-GA. The genetic algorithm used in AutoDAN-GA features a crossover rate of 0.5, a mutation rate of 0.01, and 100 optimization steps. For DrAttack and Puzzler, we use GPT-4 to construct their jailbreak prompts. For MultiJail, we selected all 315 queries in the Bengali language. For AlpacaEval, we select all 805 questions from the AlpacaEval dataset.
Metrics. We measure the defense effectiveness of SelfDefend indirectly using the attack success rate (ASR) [29] or unsafe rate [34]. Following [29, 28], ASR is measured by detecting whether the LLM response contains keywords indicating a refusal to answer, like “I can’t assist.” If such a keyword or its variant is included, it indicates that the attack is unsuccessful; otherwise, it indicates that the jailbreak is successful. We use the common practice [29] for a list of such keywords, which is available in Appendix A. GCG, JSAA, PAIR, AutoDAN, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance. The lower the ASR, the stronger the defense performance of SelfDefend against jailbreaks. Note that we also use ASR to evaluate the performance of SelfDefend under normal prompts from AlpacaEval, but a higher ASR indicates that our framework is more compatible with normal prompts (in such a case, ASR can be interpreted as answer success rate). On the other hand, the unsafe rate uses GPT-4 to determine whether the response of the target model matches the jailbreak goal, following [34]. We adopt the unsafe rate in a way similar to ASK to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.

### v2

@@ 3.1 Design Objectives
Handling all kinds of jailbreak attacks. The proposed defense should be able to handle all categories of jailbreak attacks listed in §2, including not only human-based jailbreaks but also optimization-based, generation-based, indirect, and multilingual jailbreaks.
@@
Incurring negligible delay to user prompts. The defense should not impact the user experience, causing either no delay or only a negligible one to normal user prompts.
@@
Providing explanations for potential jailbreak queries. Similar to O2, when the defense detects any query potentially related to jailbreaking, it should provide helpful explanations on why the query is considered harmful.
@@
Compatible with both open-source and closed-source LLMs. The proposed jailbreak defense approach should protect both white-box and open-source LLMs as well as black-box and closed-source LLMs.
We clarify that compared with the other three objectives, O3 is not mandatory. Nevertheless, we believe O3 is valuable for users to forensically understand potential jailbreak queries and better protect against jailbreak attacks in the future.
Figure 1: A high-level overview of the SelfDefend framework and its workflow; see §4 for more details.
@@ 3.2 Analysis of Existing Defenses
Table 1 summarizes our analysis of major LLM jailbreak defenses under the four objectives we envisioned above. They can be roughly categorized into plugin-based (the first 11 rows) and model-based (the last nine rows) mechanisms. Specifically, plugin-based defenses can be typically applied to any off-the-shelf LLMs like a plugin to enhance their safety against jailbreak attacks, while model-based defenses aim to fundamentally improve a model’s safety alignment against jailbreaks by changing the model’s internal mechanisms or conducting fine-tuning for parameter optimization. It is worth noting that some defenses may exhibit characteristics of both types, for which we do not aim to distinctly distinguish them in this paper. In the rest of this subsection, we present our analysis results from four perspectives:
@@
First, most jailbreak defenses target multiple kinds of jailbreak attacks but typically do not cover advanced indirect attacks (O1: ), while a few existing defense mechanisms are specifically designed to defend against optimization-based adversarial attacks only (O1: ). The latter includes two perplexity-based filtering approaches [4, 26] and several input perturbation-based approaches [29, 9, 59, 27]. Specifically, Alon and Kamfonas [4] proposed the first plugin-based defense in the sense that they not only leveraged perplexity values as an indicator to detect prompts with adversarial suffixes but also tuned a classifier to consider both the sequence length of prompts and their perplexity for improved filtering. Another line of GCG-specific jailbreak defenses perturbed copies of the input prompt and aggregated the output responses, with SmoothLLM [59] as a representative example.
@@
Second, the majority of plugin-based defenses inherently incur additional delays to user prompts (O2: ), while most model-based methods do not (O2: ). Since the design principles of most prior defenses are to conduct extra-round analyses of the input prompt [4, 26, 25, 88, 29], or to check the target LLM’s internal states [79, 23] and responses [25, 55, 9, 59, 27], it is thus inherently difficult for these approaches to avoid additional delay. The exceptions are most model-based defenses, which either conduct prompt tuning [96, 81, 93, 49] or optimize the parameters of the target model [89, 77, 39, 45, 91]. As a result, the tuned models behave like normal LLMs, incurring no extra delay to user prompts.
@@
Third, more than half of plugin-based defenses have the potential to provide explanations for jailbreak queries (O3: or ), whereas most model-based defenses cannot because they rely solely on LLMs’ internal mechanism tuning (O3: ). To provide explanations for potential jailbreak queries, a defense scheme needs to understand the semantics of incoming prompt queries. Hence, it is difficult for approaches that rely solely on target LLMs’ internal indicators, whether they are plugin-based [73, 80, 79, 23] or model-based [82, 91], to provide straightforward explanations to users.
@@
Fourth, most plugin-based defenses are compatible with both open-source and closed-source LLMs (O4: ), while the opposite is true for most model-based defenses (O4: ). Unless they need to monitor LLMs’ internal indicators [79, 23], plugin-based defenses, such as Self Defense [55] and IAPrompt [88], are naturally compatible with both open-source and closed-source LLMs. By contrast, model-based methods typically require white-box access to the LLMs to enable their in-depth defense. Exceptions include approaches that enhance the safety of the LLM through the integration of hand-crafted prompts [89].
@@ 4 The SelfDefend Framework
Overview. Based on this idea, we propose a generic LLM jailbreak defense framework called SelfDefend. As shown in Figure 1, SelfDefend creatively establishes a shadow stack alongside the normal stack in the LLM space to conduct checkpoint-based access control, which mimics traditional security defense concepts such as the shadow stack for defending against buffer overflow attacks [8] and the library-based checkpoint from SCLib [74]. We denote the target LLM in the normal stack as $LLM_{target}$ and the defense LLM in the shadow stack as $LLM_{defense}$. SelfDefend simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate. In SelfDefend, $LLM_{defense}$ can be instantiated from the same model as $LLM_{target}$, although in practice we suggest using a dedicated $LLM_{defense}$ that is robust and low-cost for detecting jailbreak queries.
Workflow. ① Given an incoming prompt query $P_{query}$, SelfDefend dispatches it to both $LLM_{target}$ and $LLM_{defense}$ for concurrent processing. ② $LLM_{target}$ processes $P_{query}$ as usual, whether it is a normal prompt or an adversarial prompt, but caches its token-by-token output until a checkpoint is triggered from the shadow stack. By contrast, $LLM_{defense}$ employs a tailored detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful part (via $P_{direct}$) or intention (via $P_{intent}$). ③ Once a token of “No” (indicating no issue) is output from the shadow stack, $LLM_{defense}$ is triggered to release its token-by-token response. Otherwise, when the shadow stack detects a harmful portion (prompt/intention), SelfDefend would respond with a template that refuses to answer, i.e., “I can’t fulfill your query because your [harmful portion] violated our safety policy.” It is noted that “[harmful portion]” is replaced with the recognized portion from $LLM_{defense}$.
Objective Analysis. Since the time for outputting the token “No” is very limited, the additional delay introduced by SelfDefend is negligible for normal user prompts (O2: ). Moreover, the identified harmful portion can help generate an explainable LLM response to adversarial prompts (O3: ). These unique advantages make SelfDefend the first practical jailbreak defense compared to the existing defense mechanisms explained in §3.2. Moreover, as $LLM_{defense}$ does not need to modify $LLM_{target}$ or know any of its internal details, SelfDefend can protect all kinds of target LLMs, whether they are open-source or closed-source (O4: ).
Now the effectiveness of SelfDefend depends on under which detection prompts, it can handle all kinds of jailbreak attacks (i.e., O1). For this, we proceed to conduct an empirical study in §5 to explore effective $P_{direct}$ and $P_{intent}$ and validate our earlier hypothesis that “LLMs have the capability to identify harmful portions in user queries” using these measures.
@@ 5 An Empirical Measurement
In this section, we conduct extensive measurements to empirically show that under the SelfDefend framework and its carefully designed detection prompts, existing LLMs, notably the commonly used GPT-3.5/4 models, perform acceptably. This not only validates the hypothesis assumed in §4 but also enables widespread, convenient usage of SelfDefend without the need for further tuning of a defense model. For example, various custom GPTs [1] can adopt SelfDefend to immediately enhance their defense against jailbreaks.
5.1 Initializing SelfDefend for GPT-3.5/4 and Designing Two Detection Prompts
@@
To facilitate measurements, we need to first initialize SelfDefend for GPT-3.5/4. For GPT-3.5-based experiments, both $LLM_{target}$ in the normal stack and $LLM_{defense}$ in the shadow stack are set to GPT-3.5. Likewise, for GPT-4-based experiments, both $LLM_{target}$ and $LLM_{defense}$ are set to GPT-4. The API versions of GPT-3.5 and GPT-4 are gpt-3.5-turbo-0125 and gpt-4-0125-preview, respectively. We set the hyper-parameters max_tokens, temperature, n, and top_p to 128, 0, 1, and 0, respectively. When accessing the GPT APIs, no system prompt is added.
Figure 2: The two detection prompt templates designed. “[Jailbreak/Normal Prompt]” will be replaced with the user query.
Besides initializing SelfDefend for GPT-3.5/4, we carefully design $P_{direct}$ and $P_{intent}$ for SelfDefend and evaluate their effectiveness in this section. As mentioned in §4, $P_{direct}$ and $P_{intent}$ represent two different defense schemes. $P_{direct}$ is more intuitive, aiming to directly pinpoint the harmful part from the original jailbreak prompt; otherwise, it outputs a token of “No”. $P_{intent}$, on the other hand, aims to recognize the true intent of the request text rather than directly intercepting harmful parts, given that indirect jailbreak attacks (e.g., DrAttack [33]) hide harmful instructions.
@@
Based on these two guidelines, we carefully design the prompt templates used by $P_{direct}$ and $P_{intent}$. As shown in Figure 2, $P_{direct}$ uses the sentence like “identify which part of the following sentences might violate your safety policies or have the jailbreak intentions” to directly pinpoint the harmful portion. Inspired by chain-of-thought (CoT) [71], $P_{intent}$ first summarizes the true intention of $P_{query}$ and then, further extracts any safety-violation part from the summarized intention, if present. It also includes two contrasting examples in the prompts to help $LLM_{defense}$ better understand the instructions, motivated by few-shot in-context learning [7].
Through the extensive measurements in this section, these two designed prompts are shown to be effective within the SelfDefend framework. Nevertheless, we note that the detailed wording of these two prompt contents might have alternatives, but the current form is used as a standard template for $P_{direct}$ and $P_{intent}$ throughout the paper.
5.2 Datasets and Attack Setup
Benchmarks. Based on the five categories of existing jailbreak attacks we surveyed in §2 — human-based, optimization-based, generation-based, indirect, and multilingual jailbreaks — we identify representative jailbreak attack methods in each category. We then collect four benchmark datasets, JailbreakHub [62], JailbreakBench [11], MultiJail [17], and AlpacaEval [34], from which we use their user prompts for testing SelfDefend. Table 2 lists the details of our collected benchmark datasets. Specifically, we use a set of 100 harmful instructions from JailbreakBench [11], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [99] and AutoDAN [40]), generation-based jailbreaks (PAIR [12] and TAP [47]), and indirect jailbreak attacks (DrAttack [33] and Puzzler [10]). In contrast, we directly use the original prompts from the JailbreakHub, MultiJail and AlpacaEval datasets to construct inputs for the scenarios of human-based attacks, multilingual jailbreaks and normal prompts, respectively.
Table 2: The details of our collected benchmark datasets.
@@ 5.2 Datasets and Attack Setup
JailbreakHub [62] 1000 DAN [62]
JailbreakBench [11] 100 GCG [99], AutoDAN [40]
PAIR [12], TAP [47]
@@
Table 3: The ASR results from testing LLMs against five major categories of jailbreak attacks and normal prompts.
LLMs Human-based Optimization-based Generation-based Indirect Multilingual Normal
DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 (baseline) 0.256 0.560 0.900 0.720 0.670 0.780 0.980 0.393 0.977
GPT-3.5-based Shadow Stack ($P_{direct}$) 0.982 0.720 0.960 0.770 0.840 1.000 0.980 0.879 0.977
GPT-3.5-based SelfDefend ($P_{direct}$) 0.242 0.450 0.870 0.600 0.610 0.780 0.960 0.368 0.957
GPT-3.5-based Shadow Stack ($P_{intent}$) 0.015 0.280 0.350 0.370 0.020 0.860 0.220 0.520 0.992
GPT-3.5-based SelfDefend ($P_{intent}$) 0.007 0.190 0.310 0.290 0.020 0.710 0.220 0.203 0.972
Reduction factor 97.26% 66.07% 65.55% 59.72% 97.01% 8.97% 77.55% 48.34% 0.51%
GPT-4 (baseline) 0.047 0.080 0.190 0.330 0.310 0.740 0.900 0.076 0.973
GPT-4-based Shadow Stack ($P_{direct}$) 0.004 0.010 0.010 0.110 0.100 0.050 0.270 0.142 0.968
GPT-4-based SelfDefend ($P_{direct}$) 0.002 0.000 0.010 0.100 0.080 0.040 0.260 0.012 0.946
Reduction factor 95.74% 100.00% 94.73% 69.69% 74.19% 94.59% 71.11% 84.21% 2.77%
GPT-4-based Shadow Stack ($P_{intent}$) 0.019 0.080 0.070 0.210 0.200 0.130 0.360 0.304 0.995
GPT-4-based SelfDefend ($P_{intent}$) 0.005 0.050 0.070 0.180 0.170 0.130 0.280 0.019 0.970
@@
Figure 3: The extra delay $\Delta d$ caused by SelfDefend for normal and jailbreak prompts, respectively.
Attack Setup. For DAN, we randomly select 1,000 samples as jailbreak queries from the forbidden question set equipped with jailbreak prompts [2], which is collected by JailbreakHub. For GCG, we choose its individual version and optimize the suffix on Vicuna-7b-v1.3 [13] with a batch size of 512 and 500 optimization steps. For AutoDAN, we choose its first version of the genetic algorithm, i.e., AutoDAN-GA. The genetic algorithm used in AutoDAN-GA features a crossover rate of 0.5, a mutation rate of 0.01, and 100 optimization steps. For PAIR and TAP, we select Vicuna-13b-v1.5 [13] as the attack model. We set the maximum depth, the maximum width, and the branching factor of TAP to 10, 10, and 1, respectively. The target model of PAIR and TAP is GPT-3.5/4. For DrAttack and Puzzler, we use GPT-4 to construct their jailbreak prompts. For MultiJail, we selected all 315 queries in the Bengali language. For AlpacaEval, we select all 805 questions from the AlpacaEval dataset.
Metrics. We measure the defense effectiveness of SelfDefend indirectly using the attack success rate (ASR) [99] or unsafe rate [17]. Following [99, 40], ASR is measured by detecting whether the LLM response contains keywords indicating a refusal to answer, like “I can’t assist.” If such a keyword or its variant is included, it indicates that the attack is unsuccessful; otherwise, it indicates that the jailbreak is successful. We use the common practice [99] for a list of such keywords; see Appendix B. DAN, GCG, AutoDAN, PAIR, TAP, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance. The lower the ASR, the stronger the defense performance of SelfDefend against jailbreaks. Note that we also use ASR to evaluate the performance of SelfDefend under normal prompts from AlpacaEval, but a higher ASR indicates that our framework is more compatible with normal prompts (in such cases, ASR can be interpreted as answer success rate). On the other hand, the unsafe rate uses GPT-4 to determine whether the response of the target model matches the jailbreak goal [17]. We adopt the unsafe rate in a way similar to ASR to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.